# PathDB: A system for evaluating regular path queries

### v1

@@ Abstract.
PathDB is a Java-based graph database designed for in-memory data loading and querying. By utilizing Regular Path Queries (RPQ) and a closed path algebra, PathDB processes paths through its three main components: the parser, the logical plan, and the physical plan. This modular design allows for targeted optimizations and modifications without impacting overall functionality. Benchmark experiments illustrate PathDB’s execution times and flexibility in handling dynamic and complex path queries, compared to baseline methods like Depth-First Search (DFS) and Breadth-First Search (BFS) guided by an automaton, highlighting its optimizations that contribute to its performance.
@@ 1. Introduction
In the rapidly evolving field of graph databases, several systems have emerged, each showcasing distinct strengths and capabilities. Among the most prominent players are Neo4j, Kuzu and Memgraph. While these systems support path querying, they share a common characteristic: the reliance on specific algorithms to process recursive paths queries. However, Neo4j, Kuzu, and Memgraph offer only limited support for path queries.
PathDB focuses on utilizing execution trees instead of recursive algorithms, as used by most systems. This approach allows for future optimization of these trees through techniques such as predicate pushdown and other enhancements. Its uniqueness lies in a path manipulation algebra, inspired by relational algebra, based on the concept of sets of paths. This path algebra consists of three main operators: selection, which filters paths; join, which concatenates compatible paths; and union, which merges sets of paths. It is worth noting that all these operators function exclusively on sets of paths.
PathDB’s design features three main components: the parser, the logical plan, and the physical plan. By utilizing a path query pattern inspired by the GQL standard (2024iso, ), PathDB enables users to query paths and return a set of paths as results.
The primary objective of this paper is to showcase the innovative features and capabilities of PathDB. We provide a comprehensive overview of how graph data is stored, the query evaluation workflow, the syntax and operation of its query language, and related optimizations. Additionally, we demonstrate how PathDB’s design allows for targeted optimizations and modifications without disrupting overall functionality. This includes optimizations in the logical plan, changes to the physical plan, and the ability to seamlessly transition between different storage systems.
Section 2 provides a system overview, describing the architecture and data storage. Section 3 presents a demonstration of PathDB’s interface and usage, including a comparison with base methods like DFS and BFS with an automaton. Finally, Section 4 offers conclusions and future work.
@@ 2. System overview
In this section, we provide a comprehensive overview of the principal PathDB features. We will cover how the data is stored, the architecture’s design, the operation of its query language, and related optimizations.
It is worth noting that PathDB is a Java-based system designed to load and query data in memory. It features an architecture with three main components: the parser, the logical plan, and the physical plan. By using a path query pattern based on Regular Path Queries (RPQ), PathDB enables users to query for paths and return a set of paths as results. Importantly, PathDB employs a closed path algebra, meaning that each operator takes a set of paths as input and returns a set of paths as output.
@@
The information about nodes and edges is stored in a data structure that uses a compressed sparse row with vertical partitioning (CSR VP) as its base. Essentially, a CSR is represented by two lists: the first list maintains the offsets into the second array, marking the beginning of the outgoing edges for the corresponding vertex. The end of this list is marked by the beginning of the next vertex. The second array contains the edge identifiers (Tetzel2018, ). Combining CSR with vertical partitioning (VP) allows us to leverage the characteristics of a CSR while using labels, which is beneficial for our case since Regular Path Queries make use of these labels (Tetzel2018, ).
In our implementation, we have two hashmaps: one where the key is the label and the value is a Node, and another where the key is the label and the value is a linked list of Edges that contain this label. It is worth mentioning that, so far, we have demonstrated how the graph is stored, but not how to work with paths. Since the logical and physical operators in PathDB handle sets of paths, it is necessary to transform the nodes and edges of a graph $G$ into sets of paths. To achieve this, two main functions are used: $\texttt{Paths}_{0}(G)$, which converts the set of nodes into paths of length 0, and $\texttt{Paths}_{1}(G,l)$, which transforms the edges of a graph $G$ with a specific label $l$ into a set of paths of length 1.
@@
Our path algebra (EDBT.2025.63, ) consists of four fundamental operators, which can be composed and evaluated over graph-structured data. The selection operator ($\sigma$) filters paths by applying conditions over nodes or edges. The union operator ($\cup$) merges two sets of paths, $a$ and $b$, into a single set. The path join operator ($\bowtie^{\tau}$) concatenates paths from set $a$ with compatible paths from set $b$; the join may optionally be constrained by a path semantic $\tau$, which can be one of isTrail for no repeated edges, isAcyclic for no repeated nodes, or isSimple for no repeated nodes, except the first and last node. Finally, the recursive operator ($\phi^{\tau}$) applies repeated join operations to compute transitive expansions of a path set, either until a fix point is reached or until specific criteria are satisfied. Like the join, the recursive operator can also be constrained by a path semantic $\tau$, allowing it to generate only trails, simple, or acyclic paths when required. These operators form the foundation for expressing rich navigational queries. An example that demonstrates the combined application of selection, union, join, and recursion is presented in Figure 1, shaped as an execution tree.
@@
A path query pattern is an expression inspired by the GQL standard (2024iso, ), and consists of three main components: the MATCH clause, that defines the path pattern, the WHERE clause, that defines the selection conditions, and the RETURN clause, that specifies the output.
@@
The WHERE clause consists of a set of selection condition terms used to filter the paths retrieved in the MATCH clause. Let $i\geq 1$ be a natural number, pr a property label, $v$ a value, $\oplus$ a comparison operator ($=$, $\neq$, $<$, $>$, $\leq$, $\geq$), and sVar and tVar symbolic variables representing the source and target nodes of a path pattern. A selection condition $c$ is defined recursively. A simple selection condition in PathDB can be any expression such as sVar.pr $\oplus$ v, tVar.pr $\oplus$ v, FIRST().pr $\oplus$ v, LAST().pr $\oplus$ v, NODE(i).pr $\oplus$ v, EDGE(i).pr $\oplus$ v, LABEL(sVar) $\oplus$ v, LABEL(tVar) $\oplus$ v, LABEL(FIRST()) $\oplus$ v, LABEL(LAST()) $\oplus$ v, LABEL(NODE(i)) $\oplus$ v, LABEL(EDGE(i)) $\oplus$ v, LENGTH() $\oplus$ v, as well as boolean predicates such as ISTRAIL(), ISSIMPLE(), and ISACYCLIC(). If $c_{1}$ and $c_{2}$ are selection conditions, then $(c_{1}{\small\texttt{ AND }}c_{2})$ and $(c_{1}{\small\texttt{ OR }}c_{2})$ are complex selection conditions.
Finally, the RETURN clause specifies the elements of the path to be projected in the result. Projection expressions may reference any variable defined in the MATCH clause, including path and node variables. Let $i\geq 1$ be a natural number, pr a property label, Var a path pattern variable, and sVar and tVar the symbolic identifiers of the source and target nodes of that path. A PathDB projection term can refer to node or edge attributes using expressions such as sVar.pr, tVar.pr, NODE(i).pr, EDGE(i).pr, FIRST().pr, and LAST().pr. It can also include functions that extract structural or semantic features of the path, such as LABEL(), LENGTH(), ISTRAIL(), ISSIMPLE(), and ISACYCLIC(). An optional LIMIT clause may be appended to restrict the number of returned paths.
@@
The PathDB design is structured around three primary components, reflecting its query evaluation workflow, as shown in Figure 2: the parser, the logical plan, and the physical plan.
@@
Second, the AST, specifically the regular expression, serves as the input for the logical planner component. The AST is then transformed into a logical plan tree. Essentially, we convert the regular expression and endpoint filters (represented as a tree) into multiple algebraic operators. It is worth noting that our logical plan tree is executed in a bottom-up manner, from the leaves to the root.
For example, to retrieve paths that start at a node named $Moe$ and traverse one or more repetitions of either a Knows or Likes edge followed by a HasCreator edge, a user can formulate the query:
@@
The above query is based on the graph of Figure 1, the corresponding logical plan is presented in Figure 3(a). In this logical plan, the tree initially retrieves all edges from the graph $G$ labeled with Knows, Likes and HasCreator, subsequently transforming these edges into a set of paths. If we evaluate the expression from the left-hand side of the union, a recursive operator is applied over the set of paths labeled with Likes. On the right-hand side, a recursive operator is applied over the Join of two path sets: one labeled with Likes and the other with HasCreator. It is important to note that the path semantic isTrail is enforced on the resulting paths—both for the recursion and for the join—ensuring that no edge is repeated during evaluation.
Once the recursion completes, we filter the resulting paths to retain only those whose source node has the property "Name = Moe". Finally, we return the names of the individuals in the last node of each path, limiting the output to 10 results.
As can be observed, since we are working with a tree structure where each operation processes a set of paths, various optimizations can be applied during execution, which could improve execution times. These optimizations are discussed further later on.
Finally, the output of the logical plan is used to retrieve results from the graph through the physical plan. To achieve this, each logical operator in the logical plan is translated into a corresponding physical operator. In this case, the physical plan operates on the graph stored in memory, processing each result on demand.
The design of PathDB allows for the isolation of these three components, thereby enabling optimizations or modifications to be made to each one without causing significant issues. This isolation, for instance, allows for optimizations in the logical plan, changes to the physical plan, such as moving from memory storage to disk storage or any other system, like a database, among other possible modifications.
@@
Leveraging the Path Algebra to represent logical plans for evaluating Regular Path Query Semantics (RPQS) offers considerable potential for optimization through query rewriting. The algebraic framework enables the systematic transformation of queries into more efficient forms while preserving their semantic integrity. A prominent optimization technique within this context is predicate pushdown, which involves moving constraints on the source or target nodes of paths closer to the leaf nodes in the logical plan. For instance, for the query discussed before, MATCH TRAIL p = (x)-[(Knows|Likes.HasCreator)+]->(y) WHERE x.name = "Moe" RETURN y.name LIMIT 10, Figure 3(a) illustrates the original query plan. Meanwhile, Figure 3(b) presents the optimized plan, where the selection of the source node has been pushed down through the union for each child on the left-hand side.
@@ 3. Demonstration
In this section, we demonstrate PathDB’s interface and usage. We showcase how to perform path queries, load data into memory, and interact with the system’s components. Additionally, we provide a comparison with base methods such as Depth-First Search (DFS) and Breadth-First Search (BFS) with an automaton to illustrate PathDB’s capabilities. PathDB is available on GitHub [^2] https://github.com/dbgutalca/PathDB.
@@ Abstract
Figure 5. Average Execution Time in Seconds by Query Type with Walk Semantic. Missing points indicate unsupported queries.
@@
Figure 6. Average Execution Time in Seconds by Query Type with Trail Semantic. Missing points indicate unsupported queries.
@@
Figure 7. Average Execution Time in Seconds by Query Type with Simple Semantic. Missing points indicate unsupported queries.

### v2

@@ Abstract.
PathDB is a Java-based graph database designed for in-memory data loading and querying. By utilizing Regular Path Queries (RPQ) and a closed path algebra, PathDB processes paths through its three main components: the parser, the logical plan, and the physical plan. This modular design allows for targeted optimizations and modifications without impacting overall functionality. Benchmark experiments illustrate PathDB’s execution times and flexibility in handling dynamic and complex path queries, compared to baseline methods like Depth-First Search (DFS) and Breadth-First Search (BFS) guided by an automaton, highlighting PathDB optimizations that contribute to its performance. PathDB was also evaluated against leading commercial graph systems, including Neo4j, Memgraph, and Kùzu. Benchmark experiments demonstrated PathDB’s competitive execution times and its ability to support a wide range of path query types.
@@ 1. Introduction
In the rapidly evolving field of graph databases, several systems have emerged, each showcasing distinct strengths and capabilities. Among the most prominent players are Neo4j, Kuzu and Memgraph. While these systems support path querying, they share a common characteristic: the reliance on specific algorithms to process recursive paths queries. In contrast, Neo4j, Kuzu, and Memgraph offer only limited support for path queries.
PathDB is a Java-based graph database designed for in-memory data loading and querying. By utilizing Regular Path Queries (RPQ) and a closed path algebra. Instead of relying on traditional graph traversal algorithms such as Breadth-First Search (BFS) or Depth-First Search (DFS), PathDB adopts a recursive query evaluation strategy centered around the use of the join operator. As a result, query evaluation in PathDB can be represented as execution trees, this approach allows for future optimization of these trees through techniques such as predicate pushdown and other enhancements. Its uniqueness lies in a path manipulation algebra, inspired by relational algebra, based on the concept of sets of paths. This path algebra consists of three main operators: selection, which filters paths; join, which concatenates compatible paths; and union, which merges sets of paths. It is worth noting that all these operators function exclusively on sets of paths.
PathDB’s design features three main components: the parser that perform the query validation and extraction, the logical plan responsible for converting queries into a logical plan and applying subsequent optimizations, and the physical plan which extracts the paths to be consulted using the different algebra operators.
PathDB’s query language uses a path pattern inspired by the GQL standard (2024iso, ), allowing users to query paths and retrieve sets of paths.
The primary objective of this paper is to showcase the innovative features and capabilities of PathDB. We present a comprehensive overview of its graph data storage model, configuration parameters, and query interface. Furthermore, we describe the query evaluation workflow, detailing the syntax and operation of its query language as well as related optimizations. Additionally, we demonstrate how PathDB’s design allows for targeted optimizations and operator modifications without disrupting overall functionality.
Section 2 provides a system overview, describing the architecture and data storage. Section 3 presents a demonstration of PathDB’s interface and usage, including a comparison with base methods like DFS and BFS with an automaton. Finally, Section 4 presents the conclusions and the future work.
@@ 2. System overview
In this section, we provide a comprehensive overview of the principal PathDB features. We cover how the data is stored, the architecture’s design, the query language features, and logical plan optimizations.
It is worth noting that PathDB is a Java-based system designed to load and query data in memory. It features an architecture with three main components: the parser, the logical plan, and the physical plan. By using a path query pattern based on Regular Path Queries (RPQ), PathDB enables users to query for paths and return a set of paths as results. Importantly, PathDB implements a closed path algebra, meaning that each operator takes a set of paths as input and returns a set of paths as output.
@@
The information about nodes and edges is stored in a data structure that uses a compressed sparse row with vertical partitioning (CSR VP) as its base. Essentially, a CSR (Tetzel2018, ) is represented by two lists: the first list, often called offsets, stores the index positions that indicate where each vertex’s list of outgoing edges begins in the second list, as illustrated in Figure 4. The second array, commonly referred to as columns, contains the destination vertices of all edges, flattened into a single list. Combining CSR with vertical partitioning (VP) allows us to leverage the characteristics of a CSR while using labels, which is beneficial for our case since Regular Path Queries make use of these labels. It is important to note that when the graph contains too many labels, VP becomes inefficient: it creates many small tables, leading to more joins, more reads, and worse performance.
Offsets0 2 5 6 Edge IDse0 e2 e3 e4 e5 e1 e6 e7
Figure 4. Compressed sparse row CSR example.
In our CSR based design, we define hashmaps: one where the key is the label and the value is a Node, and another where the key is an edge label and the value is a linked list of edges that contain this label. It is worth mentioning that, so far, we have shown how the graph is stored, but not how to work with paths. Since the logical and physical operators in PathDB handle sets of paths, we need to transform the nodes and edges of a graph $G$ into sets of paths. To achieve this, two main functions are used: $\texttt{Paths}_{0}(G)$, which retrieves the paths of length 0 of a graph $G$, and $\texttt{Paths}_{1}(G,l)$, which retrieve the paths of length 1 with a label $l$ from $G$.
@@
Our path algebra (EDBT.2025.63, ) defines five core operators that operate over sets of paths. The selection operator ($\sigma$) filters paths by applying conditions over nodes or edges. The union operator ($\cup$) merges two sets of paths, $a$ and $b$, into a single set. The join operator ($\bowtie^{\tau}$) concatenates paths from set $a$ with compatible paths from set $b$; the join may optionally be constrained by a path semantic $\tau$, which can be one of isTrail for no repeated edges, isAcyclic for no repeated nodes, or isSimple for no repeated nodes, except the first and last node. Finally, the recursive operator ($\phi^{\tau}$) applies repeated join operations to compute transitive expansions of a path set, either until a fix point is reached or until specific criteria are satisfied. Like the join, the recursive operator can also be constrained by a path semantic $\tau$, allowing it to generate only trails, simple, or acyclic paths when required. Finally, the projection operator allows us to extract specific components from the resulting paths and optionally limit the number of results returned. These operators form the foundation for expressing rich navigational queries. An example that demonstrates the combined application of selection, union, join, recursion, and projection is presented in Figure 1, shaped as an execution tree.
@@
A path query pattern is an expression inspired by the GQL standard (2024iso, ), and consists of three main components: the MATCH clause, that indicate the path pattern, the WHERE clause, that indicate the selection conditions, and the RETURN clause, that indicates the output terms.
@@
The WHERE clause contains a selection condition used to filter the paths retrieved in the MATCH clause. Let $i\geq 1$ be a natural number, pr a property label, $v$ a value, $\oplus$ a comparison operator ($=$, $\neq$, $<$, $>$, $\leq$, $\geq$), and sVar and tVar symbolic variables representing the source and target nodes of a path pattern. A selection condition $c$ is defined recursively. A selection condition in PathDB can be any expression such as:
  • sVar.pr $\oplus$ v
  • tVar.pr $\oplus$ v
  • FIRST().pr $\oplus$ v
  • LAST().pr $\oplus$ v
  • NODE(i).pr $\oplus$ v
  • EDGE(i).pr $\oplus$ v
  • LABEL(sVar) $\oplus$ v
  • LABEL(tVar) $\oplus$ v
  • LABEL(FIRST()) $\oplus$ v
  • LABEL(LAST()) $\oplus$ v
  • LABEL(NODE(i)) $\oplus$ v
  • LABEL(EDGE(i)) $\oplus$ v
  • LENGTH() $\oplus$ v
Additionally boolean predicates such as ISTRAIL(), ISSIMPLE(), and ISACYCLIC() can be applied to filter path results. If $c_{1}$ and $c_{2}$ are selection conditions, then $(c_{1}{\small\texttt{ AND }}c_{2})$ and $(c_{1}{\small\texttt{ OR }}c_{2})$ are complex selection conditions.
Finally, the RETURN clause is used to specify the elements of the path to be projected in the result. Projection expressions may reference any variable defined in the MATCH clause, including path and node variables. Let $i\geq 1$ be a natural number, pr a property label, Var a path pattern variable, and sVar and tVar the symbolic identifiers of the source and target nodes of that path. A PathDB projection term can refer to node or edge attributes using expressions such as sVar.pr, tVar.pr, NODE(i).pr, EDGE(i).pr, FIRST().pr, and LAST().pr. It can also include functions that extract structural or semantic features of the path, such as LABEL(), LENGTH(), ISTRAIL(), ISSIMPLE(), and ISACYCLIC(). An optional LIMIT clause may be appended to restrict the number of returned results.
@@
The PathDB design is structured around three primary components, reflecting its query evaluation workflow, as shown in Figure 2: the parser, the logical planner, and the physical planner.
@@
Second, the AST, specifically the regular expression, serves as the input for the logical planner component. The AST is then transformed into a logical plan tree. Essentially, we convert the path query pattern into multiple algebraic operators. It is worth noting that our logical plan tree is evaluated in a bottom-up manner, from the leaves to the root.
For example, to retrieve paths that start at a with a property $name=Moe$ and traverse one or more repetitions of either a Knows or Likes edge followed by a HasCreator edge, a user can formulate the query:
@@
The above query is based on the graph of Figure 1, the corresponding logical plan is presented in Figure 3(a). This logical plan specifies the operators required to evaluate the query: it first retrieves all edges from the graph $G$ labeled with Knows, Likes and HasCreator, and then transforms these edges into a set of paths. If we evaluate the expression from the left-hand side of the union, a recursive operator is applied over the set of paths labeled with Likes. On the right-hand side, a recursive operator is applied over the Join of two path sets: one labeled with Likes and the other with HasCreator. It is important to note that the path semantic isTrail is enforced on the resulting paths—both for the recursion and for the join—ensuring that no edge is repeated during evaluation.
Above the union operator, the logical plan applies a selection that retains only paths whose source node has the property "Name = Moe". Finally, a projection operator extracts the names associated with the last node of each remaining path, limiting the output to the top 10 results.
As can be observed, since we are working with a tree structure where each operation processes a set of paths, various optimizations can be applied during execution, which could improve execution times.
Finally, the physical plan implements the necessary algorithms to execute the operations specified by the logical plan, retrieving the corresponding results from the graph. To achieve this, each logical operator in the logical plan is translated into a corresponding physical operator algorithm. In this case, the physical plan operates on the graph stored in memory, processing each result on demand.
The design of PathDB allows for the isolation of these three components, thereby enabling optimizations or modifications to be made to each one without causing significant issues for the rest of the operator. This isolation, for instance, allows for optimizations in the logical plan, changes to the physical plan, such as moving from memory storage to disk storage or any other system, like a database, among other possible modifications.
@@
Leveraging the Path Algebra to represent logical plans for evaluating Regular Path Query Semantics (RPQS) offers considerable potential for optimization through query rewriting. The algebraic framework enables the systematic transformation of queries into more efficient forms while preserving their semantic integrity. A prominent optimization technique within this context is predicate pushdown (10.5555/645920.672839, ), which involves moving constraints on the source or target nodes of paths closer to the leaf nodes in the logical plan. Predicate pushdown works well in practice because it minimizes the amount of data that flows through the query plan. By applying filtering conditions as early as possible.
For instance, for the query discussed before, MATCH TRAIL p = (x)-[(Knows|Likes.HasCreator)+]->(y) WHERE x.name = "Moe" RETURN y.name LIMIT 10, Figure 3(a) illustrates the original query plan. Meanwhile, Figure 3(b) presents the optimized plan, where the selection of the source node has been pushed down through the union for each child on the left-hand side.
@@ 3. Demonstration
In this section, we demonstrate PathDB’s interface and usage. We showcase how to load data into memory, perform path queries, and interact with the system’s components. Additionally, we provide a comparison with base methods such as Depth-First Search (DFS) and Breadth-First Search (BFS) with an automaton to illustrate PathDB’s capabilities. PathDB is available on GitHub [^2] https://github.com/dbgutalca/PathDB.
@@ Abstract
Figure 6. Average Execution Time per Query Type with Walk Semantics. Missing bars indicate unsupported queries.
@@
Figure 7. Average Execution Time per Query Type with Trail Semantics. Missing bars indicate unsupported queries.
@@
Figure 8. Average Execution Time per Query Type with Simple Semantics. Missing bars indicate unsupported queries.